Are phone calls equally likely to occur any day of the week? The day of the week for each of 525 randomly selected phone calls was observed. The results are displayed in the table below. Use an αα = 0.05 significance level.

  1. Complete the rest of the table by filling in the expected frequencies:
    Frequencies of Phone Calls for Each Day of the Week
    Outcome Frequency Expected Frequency
    Sunday 80
    Monday 73   
    Tuesday 69
    Wednesday 71
    Thursday 80
    Friday 77
    Saturday 75
  2. What is the correct statistical test to use?
    Select an answer Homogeneity Independence Goodness-of-Fit Paired t-test  
  3. What are the null and alternative hypotheses?
    H0:H0:
    • Phone calls and days of the week are independent.
    • The distribution of phone calls is uniform over the days of the week.
    • The distribution of phone calls is not uniform over the days of the week.
    • Phone calls and days of the week are dependent.







    H1:H1:
    • The distribution of phone calls is uniform over the days of the week.
    • Phone calls and days of the week are dependent.
    • Phone calls and days of the week are independent.
    • The distribution of phone calls is not uniform over the days of the week.
  4. The degrees of freedom =  




  5. The test-statistic for this data =  (Please show your answer to three decimal places.)




  6. The p-value for this sample = (Please show your answer to four decimal places.)




  7. The p-value is Select an answer greater than less than (or equal to)  αα




  8. Based on this, we should Select an answer fail to reject the null reject the null accept the null  
  9. Thus, the final conclusion is...
    • There is sufficient evidence to conclude that phone calls and days of the week are dependent.
    • There is insufficient evidence to conclude that phone calls and days of the week are dependent.
    • There is insufficient evidence to conclude that the distribution of phone calls is not uniform over the days of the week.
    • There is sufficient evidence to conclude that the distribution of phone calls is not uniform over the days of the week.
    • There is sufficient evidence to conclude that the distribution of phone calls is uniform over the days of the week.
Answer #1

The correct statistical test is Goodness of fit test.

H0: the distribution of phone calls is uniform over the days of week.

H1: the distribution of phone calls is not uniform over the days of week.

Categories Observed Expected Sun 80 525*0.142857=75 Mon 73 525*0.142857=75 Tue 69 525*0.142857=75 Wed 71 525*0.142857=75 Thur(1) Null and Alternative Hypotheses The following null and alternative hypotheses need to be tested: = 0.142857, P3 = 0.14285(3) Test Statistics The Chi-Squared statistic is computed as follows: n x (O; – E;) = Σ = = 0.333 + 0.053 + 0.48 +0.213 + 0.3(5) Conclusion It is concluded that the null hypothesis Ho is not rejected. Therefore, there is NOT enough evidence to claim

Ans: there is sufficient evidence to conclude that the distribution of phone calls is uniform over the days of the week.

I have given my best to solve your problem. Please like the answer if you are satisfied with it. ?